# Images from disk and spherical accretions of Bardeen black hole surrounded by perfect fluid dark matter

### v1

@@ Acknowledgements.
This work is partly supported by Natural Science Foundation of China under Grants No.12375054. Note added: When completing this article, we became aware of a related independent work by arXiv:2512.00824 Feng:2025ljz. While both studies investigate the images of Bardeen black holes surrounded by perfect fluid dark matter, our work distinguishes itself in the following key aspects: 1. We used observational data of the shadow angle diameter of black holes M87 and SgrA* to constrain the dark matter parameters; 2. More importantly, we additionally discussed images of black holes illuminated by a spherical accretion model, focusing on the influence of the accretion model on the images and investigating the relationship between the images and the accretion model. In addition, our article and reference Feng:2025ljz set different dark matter parameters, but both found that dark matter increases the inner shadow of the black hole and reduces the observation intensity, and the conclusions remained consistent.

### v2

@@ Acknowledgements.
This work is partly supported by the Natural Science Foundation of China under Grant No.12375054. The authors thank Professor Xiao-Mei Kuang and Dr. Xi-Jing Wang for their valuable help and guidance. Note added: When we finished this article, we noticed a related independent work by arXiv:2512.00824 Feng:2025ljz. Although both studies investigate the images of Bardeen black holes surrounded by PFDM, our work distinguishes itself in the following key aspects: 1. By investigating the fundamental characteristic quantities of black holes, we discovered a competing effect between dark matter parameter $\left|\alpha\right|$ and magnetic charge parameter $g$; 2. More importantly, we additionally discussed images of black holes illuminated by a spherical accretion model and investigated the relationship between the images and the accretion model. In addition, our article and reference Feng:2025ljz set different dark matter parameters, but both found that dark matter increases the inner shadow of the black hole and reduces the observation intensity, and the conclusions remained consistent.